Data Table Output level Operating assets Operating asset turnover Return on operating assets Degree of operating leverage Interest expense Tax rate (Click on the icon located on the top-right corner of the data table above in order to copy its contents into a spreadsheet.) 170,000 units $5,200,000 20 times 45% 8 times $700,000 39% Print DoneMini Case. Chapter 12 #2 (similar to) Question Help * Camping USA Inc. has beenoerating for only 2 years in the outskirts of Albuquerque, New Mexico, and is a new manufacturer of a top-of-the-line camping tent. You are starting an internship as assistant to the chief financial officer of the company, and the owner and CEO, Tom Charles, has decided that this is the right time to know more about the business and financial risks his company must deal with. For this, the CFO has asked you to prepare an analysis to support him in his next meeting with Tom Charles a week from today. To make the required calculations, you have put together the following data regarding the cost structure of the company in the popup window, BEB The CFO has instructed you to first determine the break-even point in units of output for the company. He requires that you prepare supporting documents that demonstrate how you arrived at your conclusion and can facilitate his review of your work. Accordingly, you are required to have the information needed to prepare an analytical income statement for the company to be presented to the CFO. In a format that is acceptable for a meeting discussion with the CEO, you also need to prepare answers to the following questions: a. What is the firms break-even point in sales dollars? b. If sales should increase by 20 percent, by what percentage would EBT (earnings before taxes) and net income increase? c. Prepare another income statement, this time to verify the calculations from part (b)
Answer #1

In the give question, it is important to find prepare the Income statement at first, by using all the given information.

We can calculate the total Sales amount by using the Operating Assets turnover ratio (20 times) and the Operating Assets amount ($5,200,000).

Operating Assets Turnover = Net Sales/Operating Assets

Thus, Net Sales = 20*($5,200,000) = $104,000,000.

Similarly, we can calculate the Net Income by using the Return on Operating Assets ratio (45%) and the Operating Assets amount ($5,200,000).

Return on Operating Assets = Net Income/ Operating Assets

Thus, Net Income = 45%*($5,200,000) = $2,340,000.

Using the Net Income amount, we can proceed to calculate the EBT (Earnings before Tax), by using the tax rate given as 39%. This EBT amount turns out to be $3,836,065.57

EBIT (Earnings before Interest and Tax) can be calculated by adding the Interest amount ($700,000) to EBT ($3,836,065.57). The EBIT amount turns out to be $4,536,065.57.

We can calculate the (Sales-Total Variable Cost) value, using the Degree of Operating leverage and EBIT amount.

Degree of Operating Leverage = (Sales-TVC)/EBIT

Thus, (Sales-TVC) = $36,288,524.56

Income Statement is shown below:

Particulars Amount in $
Sales $104,000,000.00
(-) Variable Cost ($67,711,475.44)
$36,288,524.56
(-) Fixed Cost ($31,752,458.99)
EBIT $4,536,065.57
(-) Interest ($700,000.00)
EBT $3,836,065.57
(-) Tax ($1,496,065.57)
EAT $2,340,000.00

(a) Firm's Break-even point in Sales dollars

Break-even Quantity = (Fixed Operating Costs+Fixed Financing costs)/(Price - Variable cost per unit)

Price per unit = Sales/ Output level = $611.76

Variable cost per unit = Variable cost/ Output level = $398.30

Break Even Quantity = ($31,752,458.99 + $700,000.00) / ($611.76 - $398.30) = 152,031 units.

Therefore Break even Sales = (152,031*$611.76) = $93,006,484.56.

(b)

A firm's degree of operating leverage (DOL) can also be defined as the percentage change in operating income (EBIT) that results from a given percentage change in sales.

DOL = (% change in EBIT)/(% change in Sales)

Therefore,

% change in EBIT = DOL * (% change in Sales)

% change in EBIT = 8* (20%) = 160%

New EBIT = $4,536,065.57*(1+160%) = $11,793,770.48.

New EBT = New EBIT - Interest = $11,793,770.48 - $700,000 = $11,093,770.48.

% change in EBT = ($11,093,770.48-$3,836,065.57)/($3,836,065.57) = 189%

This is similar to 160%*(Degree of financial leverage) = 160%(1.18) = 188.8%

DFL (Degree of Financial Leverage) can also be defined as the percentage change in Net Income that results from a given percentage changes in EBIT.

Thus, % change in Net Income = 1.18*(160%) = 188.8%

(c)

The revised Income Statement is shown below:

Particulars Amount in $
Sales $124,800,000.00
(-) Variable Cost ($81,253,825.03)
$43,546,174.97
(-) Fixed Cost ($31,752,458.99)
EBIT $11,793,715.98
(-) Interest ($700,000.00)
EBT $11,093,715.98
(-) Tax ($4,326,549.23)
EAT $6,767,166.75